# Continuous-variable quantum key distribution system: A review and perspective

### v2

@@ III.1 Quantum random number generator
Based on different security levels, the QRNG can be divided into the practical type Jennewein et al. (2000); Stefanov et al. (2000); Dynes et al. (2008); Wayne et al. (2009); Gabriel et al. (2010); Symul, Assad, and Lam (2011), the semi-device independent type Nie et al. (2016); Cao, Zhou, and Ma (2015); Avesani et al. (2018); Marangon, Vallone, and Villoresi (2017); Cao et al. (2016); Brask et al. (2017); Van Himbeeck et al. (2017); Xu et al. (2019); Tebyanian et al. (2021); Avesani et al. (2021) and the device independent type Liu et al. (2018b, 2021b); Li et al. (2021a); Zhang et al. (2020d). The QRNG can also be divided into the discrete-variable and continuous-variable types corresponding to different types of entropy source. Normally, the continuous-variable QRNG can support high generation rate with simple structure, where the randomness can come from the ASE noise Williams et al. (2010); Li et al. (2011); Wei et al. (2011); Argyris et al. (2012); Li et al. (2014b); Zhou, Yuan, and Ma (2015); Yang et al. (2020), the phase noise Qi et al. (2010a); Guo et al. (2010); Xu et al. (2012); Abellán et al. (2014); Yuan et al. (2014); Nie et al. (2015); Zhang et al. (2016); Abellan et al. (2016); Liu et al. (2016a); Yang et al. (2016); Sun and Xu (2017); Álvarez et al. (2020); Raffaelli et al. (2018a); Imran et al. (2021); Roger et al. (2019); Chrysostomidis et al. (2023) and the vacuum noise Jofre et al. (2011); Haw et al. (2015); Raffaelli et al. (2018b); Gehring et al. (2021); Bai et al. (2021); Bruynsteen et al. (2023). The QRNG based on vacuum noise can achieve high-speed and simple structure, using only a laser source and a homodyne detector Gehring et al. (2021); Bai et al. (2021); Raffaelli et al. (2018b). The state of the art QRNG based on the vacuum noise can achieve a generation rate of 100 Gbps using photonic integrated circuits Bruynsteen et al. (2021a), where the speed, size and scalability can well satisfy the need of a CV-QKD system Bruynsteen et al. (2023). Here we summarize the most commonly used QRNG implementations as shown in Table 6.
@@ III.2.2 Modulation
For practical implementations, we remark that these two methods require different acquisition Jouguet et al. (2012b). Besides that, the working parameters of the modulators should be accurately adjusted, since the imperfect modulation may increase the excess noise and open security loopholes Liu et al. (2017a).
@@
In the earliest experimental demonstrations, the quantum signals and the LOs are transmitted separately Grosshans et al. (2003a). Later, in order to avoid any polarization and phase drifts that may occur between the signal and LO over long-distance fiber transmissions, time-division multiplexing scheme was proposed to co-transmit the quantum signals and the LOs in the same fiber Lodewyck et al. (2007). In another literature, B. Qi et al. also co-transmitted the quantum signals with the LOs, but adopted a scheme combining polarization and frequency-division multiplexing Qi et al. (2007). In later experimental demonstration, the co-transmission schemes basically adopt the combination scheme of time multiplexing and polarization multiplexing Jouguet et al. (2013); Huang et al. (2016b); Li et al. (2017a); Zhang et al. (2019a, 2020a). In this way, the leakage from the LOs to the quantum signals can be maximally suppressed, while maintaining a relatively simple implementation structure.
@@ III.3.2 De-modulation
Since the quantum signal is transmitted in single-mode fiber while the receiver is a polarization-maintaining system, the compensation of polarization is crucial for reducing the loss of quantum signals. More importantly, it can suppress the excess noise caused by the crosstalk of quantum and classical auxiliary signals in a polarization-multiplexing system. For proof-of-principle experiments, the polarization compensation is realized by a manual polarization controller, while dynamic polarization controller with feedback systems is adopted in long-term or field tests Zhang et al. (2019a).
@@
Clock synchronization responses for the alignment of the sampling points at the transmitter and the receiver site. If the sampling points of the receiver are mismatched to those of the transmitter, the correlation between the modulation data and the detection data will be reduced. As for clock synchronization, there are many hardware layer solution, such as transmitting in a cable Mo et al. (2005), in a paired fiber da Silva and von der Weid (2009), multiplexing with quantum signal in one fiber by wave-length division multiplex (WDM) Beveratos et al. (2002), and distilling the synchronization signal from local oscillator. Each of the above options has its advantages and disadvantages. The electric cable is not only expensive and also instable for distant transmission. Transmitting clock signal and quantum signal in the same fiber by WDM can compensate the difference of signal arrival time but may introduce excess noise generated by WDM crosstalk. Distilling the synchronization signal from local oscillator need divide the local oscillator beam. The division of local oscillator power will decrease the maximal transmission distance or require higher gain of homodyne detection to compensate. Transmitting clock signal in another independent fiber will generate the fluctuation between quantum and clock signal.
@@ III.3.3 Detection
The bandwidth, detection efficiency, electronic noise and quantum to classical noise ratio (QCNR) are the key parameters of homodyne detectors. The bandwidth of the receiver directly affects the settings of the system in frequency domain, including the multiplexing and processing of the quantum and phase reference signals. The detection efficiency is another crucial parameter of the practical homodyne detector, corresponding to the detector parameter $\eta$ in Eq. (28). It is limited by the photodiodes, and the balance of the two arms of the homodyne detectors. The electronic noise of the practical homodyne detector is vital in the CV-QKD systems, which corresponds to $\nu_{ele}$ in Eq. (28). Since the power of the quantum signal is extremely small, a high electronic noise can significantly reduce the SNR, therefore resulting in a low-quality detection. The last parameter that are normally used to describe the homodyne detector is the quantum to classical noise ratio (QCNR). The QCNR demonstrate how good the weak quantum signals can be amplified while suppressing the electronic noise. The general requirement for the QCNR is that it should be at least 10 dB. We also list some reported homodyne detectors with their parameters in Table. 7, common-mode rejection ratio is introduced to reflect the balance of the detector.
Figure 14: (a) The chip-based homodyne detector for beyond 20 GHz shot-noise-limited measurements. From C. Bruynsteen et al. Bruynsteen et al. (2021b). (b) The chip-based homodyne detector for 9 GHz measurement of squeezed light. From J. Tasker et al. Tasker et al. (2021). (a) Reproduced with permission from Optica 8, 1146 (2021). Copyright 2021 Optica Publishing Group. (b) Reproduced with permission from Nat. Photonics 15, 11 (2021). Copyright 2021 Springer Nature Limited.
@@ III.4 Shot noise unit calibration
Here, $\hat{x}_{s}$ $\hat{x}_{v}$ and $X_{ele}$ are the quadrature information of quantum signal, vacuum state (shot noise), and electronic noise. $X_{LO}$ represents the effect of local oscillator, and $A$ is the amplification coefficient of circuits. Therefore, the SNU calibrated with the method we mentioned before is $(AX_{LO})^{2}$. In this way, after normalization we can get
@@
One-time Zhang et al. (2020e, 2019a) Real time
@@
Figure 19: The state-of-the-art digital signal processing routines of different local LO CV-QKD systems. (a) The routine of a Gaussian modulated system with polarization and frequency division multiplexing of quantum and pilot signals Wang et al. (2022b). (b) The routine of a discrete modulated dual-polarization system where the polarization compensation is finished digitally Roumestan et al. (2022). (c) The routine of a Gaussian modulated system with frequency division multiplexing of quantum and pilot signals. Machine learning (ML) is used for high-quality phase compensation with a wide range of pilot signal-to-noise ratios Chin et al. (2021); Hajomer et al. (2023b). (a) H. Wang et al., Commun. Phys., 5, 162, 2022; licensed under a Creative Commons Attribution ($C$ BY) license. (b) F. Roumestan et al., arXiv, 2207.11702, 2022; licensed under a Creative Commons Attribution ($C$ BY) license. (c) H. Chin et al., npj Quantum Inf., 7, 20, 2021; licensed under a Creative Commons Attribution ($C$ BY) license.
@@
This means the imperfect detector scheme in security analysis can be built with two beam splitters, one represents the detection efficiency, the other one represents the electronic noise, as shown in Fig. 17 (b). The beamsplitters represent the electronic noise and the detection efficiency can be can be exchanged in order. Note that, since electronic noise is not estimated in one-time SNU calibration, the output mode of the beamsplitter representing the electronic noise cannot be seen as trusted. This means the loss introduced by electronic noise is untrusted, which is the reason that the performance of the protocol with one-time SNU calibration is slightly lower than that using the traditional calibration method Zhang et al. (2020e).
@@ III.5 Digital signal processing
For the receiver, the frequency division multiplexed quantum and pilot signal are firstly separated from the output of the detection data, then the quantum signals are down converted to the baseband corresponding to the frequency shift at the transmitter. Secondly, clock synchronization is processed, which is a digital alternative to the hardware solution mentioned before. Subsequently, the static equalization algorithms are performed to compensate the impairments, including the IQ imbalance compensation and S21 compensation, and to recover the signal pulse with matched RRC filter. Down-sampling at the best sampling point is then performed to achieve the information with the best SNR. Then, the dynamic equalization is processed to compensate the phase and polarization mismatch, normally using Multiple-Input Multiple-Output (MIMO) algorithms Marie and Alléaume (2017); Corvaja (2017); Qi and Lim (2018); Zou, Mao, and Chen (2019); Wang et al. (2019c); Li et al. (2019a); Xing et al. (2022). Finally, frame synchronization is performed to align the beginning of the modulation and detection data. The final output of the DSP is treated as the raw data, which is then sent to the postprocessing process.
@@ III.6.3 Information reconciliation
Here, $H(X)$ is the Shannon information of the target bit string, leakleak represents the information leakage during the public transmission (usually the syndrome), and $I_{AB}$ is the mutual information between Alice and Bob. The efficiency of information reconciliation plays a crucial role in the final system’s secret key rate and the maximal transmission distance, which is seriously affected by the reconciliation strategy.
@@
In a long distance system, where the SNR is low, LDPC code is usually used Jiang et al. (2018); Zhang et al. (2020f); Guo et al. (2020); Yang et al. (2021); Xie et al. (2022); Sun and Liang (2023). The graphical representation of typical multi-edge-type LDPC code is shown in Fig. 23. Towards the practical application, the rate adaptive error correction is proposed to support the system with fading SNR Jiang et al. (2017); Zhou et al. (2021); Cao et al. (2023). In addition, to raise the utilization rate of the raw data, exchanging the order of parameter estimation and reconciliation is proposed. One can process the parameter estimation after the error correction. In this way, the abandoned data for parameter estimation can be reduced.
@@ III.6.4 Privacy amplification
The elements on the main diagonal of the Toplitz matrix are equal, and the elements on the lines parallel to the main diagonal are also equal. In this way, a Toeplitz matrix can be easily constructed. With the calculated secret key rate after parameter estimation, the length of the secret key bits can be distilled is known. Assuming the length of K’is L’ and the length of K is L, G can be a L $\times$ L’ Toeplitz matrix.
@@ IV.1.2 Long distance achievements
Figure 26: The in-line LO CV-QKD systems with long distance. (a) 80 km CV-QKD system from P. Jouguet et, al. Jouguet et al. (2013). (b) 100 km CV-QKD system from D. Huang et al. Huang et al. (2016b). (c) 202.81 km CV-QKD system from Y. Zhang et al. Zhang et al. (2020a). (a) Reproduced with permission from Nat. Photonics 7, 378-381 (2013). Copyright 2013 Macmillan Publishers Limited. (b) D. Huang, Sci. Rep., 6, 1-9, 2016; licensed under a Creative Commons Attribution ($C$ BY) license. (c) Reproduced with permission from Phys. Rev. Lett. 125, 010502 (2020). Copyright 2020 American Physical Society.
@@
The key modules of an in-line LO CV-QKD system are the multiplexing module and the reconciliation module. It can be seen that the development of the reconciliation directly supports the long-distance transmission, where the long transmission distance always combines with high reconciliation efficiency. The multiplexing scheme finally evolves into the time and polarization division multiplexing, since this is the simplest method to achieve sufficient isolation between quantum signal and LO.
Figure 27: The field tests of in-line LO CV-QKD systems. (a) The first field test of CV-QKD system, from S. Fossier et al. Fossier et al. (2009b). (b) The field test of using CV-QKD for classical symmetric encryption, from P. Jouguet et al. Jouguet et al. (2012a). (c) Field demonstration of a CV-QKD network, from D. Huang et al. Huang et al. (2016a). (d) The longest field test of CV-QKD with 50 km commercial fiber, from Y. Zhang et al. Zhang et al. (2019a). (a) Reproduced with permission from New J. Phys. 11, 045023 (2009). Copyright 2009 IOP Publishing Ltd. (b) Reproduced with permission from Opt. Express 20, 14030 (2012). Copyright 2012 Optical Society of America. (c) Reproduced with permission from Opt. Lett. 41, 3511 (2016). Copyright 2016 Optical Society of America. (d) Reproduced with permission from Quantum Sci. Technol. 4, 035006 (2019). Copyright 2019 IOP Publishing Ltd.
@@
Besides the reconciliation and multiplexing, the modulation, monitoring, sampling and compensation techniques are also well developed. The generation of light pulses with high extinction ratio is demonstrated in 2015 Wang et al. (2015c), where the extinction ratio overpasses 80 dB by using a double cascaded MZI modulation. Later, the imperfect quantum state preparation is theoretically and experimentally investigated in 2017, which demonstrates that the incorrect calibration of the working parameters for the amplitude modulator and phase modulator can lead to a significant increase of the excess noise and misestimate of the channel loss. Schemes for calibrating the working parameters of the modulators are proposed and demonstrated to solve the imperfect state preparation issue. An LO monitoring scheme is proposed to enhance the practical security of a CV-QKD system Liu et al. (2017b). The LO is monitored by the balanced photodiode, and the SNU is real-time calibrated. The excess noise in the system and its impact on the performance is also investigated Wang et al. (2019e). Finite sampling bandwidth of the analog-to-digital (AD) converter may lead to inaccurate results of pulse peak sampling, which is solved by a dynamic delay adjusting module and a statistical power feedback-control algorithm Li et al. (2016d).

### v3

@@ III.1 Quantum random number generator
Based on different security levels, the QRNG can be divided into the practical type Jennewein et al. (2000); Stefanov et al. (2000); Dynes et al. (2008); Wayne et al. (2009); Gabriel et al. (2010); Symul, Assad, and Lam (2011), the semi-device independent type Nie et al. (2016); Cao, Zhou, and Ma (2015); Avesani et al. (2018); Marangon, Vallone, and Villoresi (2017); Cao et al. (2016); Brask et al. (2017); Van Himbeeck et al. (2017); Xu et al. (2019); Tebyanian et al. (2021); Avesani et al. (2021) and the device independent type Liu et al. (2018b, 2021b); Li et al. (2021a); Zhang et al. (2020d). The QRNG can also be divided into the discrete-variable and continuous-variable types corresponding to different types of entropy source. Normally, the continuous-variable QRNG can support high generation rate with simple structure, where the randomness can come from the amplifier spontaneous emission (ASE) noise Williams et al. (2010); Li et al. (2011); Wei et al. (2011); Argyris et al. (2012); Li et al. (2014b); Zhou, Yuan, and Ma (2015); Yang et al. (2020), the phase noise Qi et al. (2010a); Guo et al. (2010); Xu et al. (2012); Abellán et al. (2014); Yuan et al. (2014); Nie et al. (2015); Zhang et al. (2016); Abellan et al. (2016); Liu et al. (2016a); Yang et al. (2016); Sun and Xu (2017); Álvarez et al. (2020); Raffaelli et al. (2018a); Imran et al. (2021); Roger et al. (2019); Chrysostomidis et al. (2023) and the vacuum noise Jofre et al. (2011); Haw et al. (2015); Raffaelli et al. (2018b); Gehring et al. (2021); Bai et al. (2021); Bruynsteen et al. (2023). The QRNG based on vacuum noise can achieve high-speed and simple structure, using only a laser source and a homodyne detector Gehring et al. (2021); Bai et al. (2021); Raffaelli et al. (2018b). The state of the art QRNG based on the vacuum noise can achieve a generation rate of 100 Gbps using photonic integrated circuits Bruynsteen et al. (2021a), where the speed, size and scalability can well satisfy the need of a CV-QKD system Bruynsteen et al. (2023). Here we summarize the most commonly used QRNG implementations as shown in Table 6.
@@ III.2.2 Modulation
For practical implementations, we remark that these two methods require different acquisitions Jouguet et al. (2012b). Besides that, the working parameters of the modulators should be accurately adjusted, since the imperfect modulation may increase the excess noise and open security loopholes Liu et al. (2017a).
@@
In the earliest experimental demonstrations, the quantum signals and the LOs are transmitted separately Grosshans et al. (2003a). Later, in order to avoid any polarization and phase drifts that may occur between the signal and LO over long-distance fiber transmissions, time-division multiplexing scheme was proposed to co-transmit the quantum signals and the LOs in the same fiber Lodewyck et al. (2007). In another literature, B. Qi et al. also co-transmitted the quantum signals with the LOs, but adopted a scheme combining polarization and frequency-division multiplexing Qi et al. (2007). In later experimental demonstration, the co-transmission schemes basically adopt the combination scheme of time-division multiplexing and polarization multiplexing Jouguet et al. (2013); Huang et al. (2016b); Li et al. (2017a); Zhang et al. (2019a, 2020a). In this way, the leakage from the LOs to the quantum signals can be maximally suppressed, while maintaining a relatively simple implementation structure.
@@ III.3.2 De-modulation
Since the quantum signal is transmitted in single-mode fiber while the receiver is a polarization-maintaining system, the compensation of polarization is crucial for reducing the loss of quantum signals. More importantly, it can suppress the excess noise caused by the crosstalk of quantum and classical auxiliary signals in a polarization-multiplexing system. For proof-of-principle experiments, the polarization compensation is realized by a manual polarization controller, while dynamic polarization controller with feedback systems is adopted in long-term or field tests Zhang et al. (2019a). For simpler deployments, the digital polarization compensation is a crucial technique, where the recently proposed polarization tracking algorithm specially designed for CV-QKD system has shown the possibility of realizing a simple and effective digital polarization compensation for extremely weak quantum signals Pan et al. (2023a).
@@
Clock synchronization responses for the alignment of the sampling points at the transmitter and the receiver site. If the sampling points of the receiver are mismatched to those of the transmitter, the correlation between the modulation data and the detection data will be reduced. As for clock synchronization, there are many hardware layer solutions, such as transmitting in a cable Mo et al. (2005), in a paired fiber da Silva and von der Weid (2009), multiplexing with quantum signal in one fiber by wave-length division multiplex (WDM) Beveratos et al. (2002), and distilling the synchronization signal from LO. Each of the above options has its advantages and disadvantages. The electric cable is not only expensive and also instable for distant transmission. Transmitting clock signal and quantum signal in the same fiber by WDM can compensate the difference of signal arrival time but may introduce excess noise generated by WDM crosstalk. Distilling the synchronization signal from local oscillator need divide the local oscillator beam. The division of LO power will decrease the maximal transmission distance or require higher gain of homodyne detection to compensate. Transmitting clock signal in another independent fiber will generate the fluctuation between quantum and clock signal.
@@ III.3.3 Detection
The bandwidth, detection efficiency, electronic noise and quantum to classical noise ratio (QCNR) are the key parameters of homodyne detectors. The bandwidth of the receiver directly affects the settings of the system in frequency domain, including the multiplexing and processing of the quantum and phase reference signals. The detection efficiency is another crucial parameter of the practical homodyne detector, corresponding to the detector parameter $\eta$ in Eq. (28). It is limited by the photodiodes, and the balance of the two arms of the homodyne detectors. The electronic noise of the practical homodyne detector is vital in the CV-QKD systems, which corresponds to $\nu_{ele}$ in Eq. (28). Since the power of the quantum signal is extremely small, a high electronic noise can significantly reduce the SNR, therefore resulting in a low-quality detection. The last parameter that are normally used to describe the homodyne detector is the QCNR. The QCNR demonstrate how good the weak quantum signals can be amplified while suppressing the electronic noise. The general requirement for the QCNR is that it should be at least 10 dB. We also list some reported homodyne detectors with their parameters in Table. 7, common-mode rejection ratio is introduced to reflect the balance of the detector.
Figure 14: (a) The chip-based homodyne detector for beyond 20 GHz shot-noise-limited measurements. From C. Bruynsteen et al. Bruynsteen et al. (2021b). (b) The chip-based homodyne detector for 9 GHz measurement of squeezed light. From J. Tasker et al. Tasker et al. (2021). (a) Reproduced with permission from Optica 8, 1146 (2021). Copyright 2021 Optica Publishing Group. (b) Reproduced with permission from Nat. Photonics 15, 11 (2021). Copyright 2021 Springer Nature.
@@ III.4 Shot noise unit calibration
Here, $\hat{x}_{s}$ $\hat{x}_{v}$ and $X_{ele}$ are the quadrature information of quantum signal, vacuum state (shot noise), and electronic noise. $X_{LO}$ represents the effect of LO, and $A$ is the amplification coefficient of circuits. Therefore, the SNU calibrated with the method we mentioned before is $(AX_{LO})^{2}$. In this way, after normalization we can get
@@
One-time Zhang et al. (2020e, 2019a) Real-time
@@
Figure 19: The state-of-the-art digital signal processing routines of different local LO CV-QKD systems. (a) The routine of a discrete modulated system with polarization and frequency division multiplexing of quantum and pilot signals Wang et al. (2022b). (b) The routine of a discrete modulated dual-polarization system where the polarization compensation is finished digitally Roumestan et al. (2022). (c) The routine of a Gaussian modulated system with frequency division multiplexing of quantum and pilot signals. Machine learning (ML) is used for high-quality phase compensation with a wide range of pilot signal-to-noise ratios Chin et al. (2021); Hajomer et al. (2023b). (a) H. Wang et al., Commun. Phys., 5, 162, 2022; licensed under a Creative Commons Attribution ($C$ BY) license. (b) F. Roumestan et al., arXiv, 2207.11702, 2022; licensed under a Creative Commons Attribution ($C$ BY) license. (c) H. Chin et al., npj Quantum Inf., 7, 20, 2021; licensed under a Creative Commons Attribution ($C$ BY) license.
@@
This means the imperfect detector scheme in security analysis can be built with two beamsplitters, one represents the detection efficiency, the other one represents the electronic noise, as shown in Fig. 17 (b). The beamsplitters represent the electronic noise and the detection efficiency can be exchanged in order. Note that, since electronic noise is not estimated in one-time SNU calibration, the output mode of the beamsplitter representing the electronic noise cannot be seen as trusted. This means the loss introduced by electronic noise is untrusted, which is the reason that the performance of the protocol with one-time SNU calibration is slightly lower than that using the traditional calibration method Zhang et al. (2020e).
@@ III.5 Digital signal processing
For the receiver, the frequency-division multiplexed quantum and pilot signal are firstly separated from the output of the detection data, then the quantum signals are down converted to the baseband corresponding to the frequency shift at the transmitter. Secondly, clock synchronization is processed, which is a digital alternative to the hardware solution mentioned before. Subsequently, the static equalization algorithms are performed to compensate the impairments, including the IQ imbalance compensation and S21 compensation, and to recover the signal pulse with matched RRC filter. Down-sampling at the best sampling point is then performed to achieve the information with the best SNR. Then, the dynamic equalization is processed to compensate the phase and polarization mismatch, normally using Multiple-Input Multiple-Output (MIMO) algorithms Marie and Alléaume (2017); Corvaja (2017); Qi and Lim (2018); Zou, Mao, and Chen (2019); Wang et al. (2019c); Li et al. (2019a); Xing et al. (2022). Finally, frame synchronization is performed to align the beginning of the modulation and detection data. The final output of the DSP is treated as the raw data, which is then sent to the postprocessing process.
@@ III.6.3 Information reconciliation
Here, $H(X)$ is the Shannon information of the target bit string, leakleak represents the information leakage during the public transmission (usually the syndrome), and $I_{AB}$ is the mutual information between Alice and Bob. The efficiency of information reconciliation plays a crucial role in the final system’s secret key rate and the maximal transmission distance, which are seriously affected by the reconciliation strategy. We remark that, the purpose of the reconciliation in a practical system is to optimize the whole system, rather than simply chasing the highest efficiency. It is necessary to balance all resources of the system under a given hardware condition to maximize the transmission distance or secret key rate. The comprehensive optimization method proposed by L. Ma et al. which considers the simultaneous optimization of the modulation variance and error correction matrix achieves a significant improvement of the system performance by at least 24 % compared with the previously used optimization methods Ma et al. (2023).
@@
In a long distance system where the SNR is low, LDPC code is usually used Jiang et al. (2018); Zhang et al. (2020f); Guo et al. (2020); Yang et al. (2021); Xie et al. (2022); Sun and Liang (2023); Zhou et al. (2023). The graphical representation of a typical multi-edge-type LDPC code is shown in Fig. 23. The state-of-the-art decoding throughput can reach 1.44 Gbps and 0.78 Gbps for the code rates 0.2 and 0.1, which can support the real-time secret key generation with the speed of 71.89 Mbps and 9.97 Mbp in 25 km and 50 km Zhou et al. (2022, 2023). Towards the practical application, the rate adaptive error correction is proposed to support the system with fading SNR Jiang et al. (2017); Zhou et al. (2021); Cao et al. (2023); Yang et al. (2024). In addition, to raise the utilization rate of the raw data, exchanging the order of parameter estimation and reconciliation is proposed. One can process the parameter estimation after the error correction. In this way, the abandoned data for parameter estimation can be reduced.
@@ III.6.4 Privacy amplification
The elements on the main diagonal of the Toeplitz matrix are equal, and the elements on the lines parallel to the main diagonal are also equal. In this way, a Toeplitz matrix can be easily constructed. With the calculated secret key rate after parameter estimation, the length of the secret key bits can be distilled is known. Assuming the length of K’is L’ and the length of K is L, G can be a L $\times$ L’ Toeplitz matrix.
@@ IV.1.2 Long distance achievements
Figure 26: The in-line LO CV-QKD systems with long distance. (a) 80 km CV-QKD system from P. Jouguet et, al. Jouguet et al. (2013). (b) 100 km CV-QKD system from D. Huang et al. Huang et al. (2016b). (c) 202.81 km CV-QKD system from Y. Zhang et al. Zhang et al. (2020a). (a) Reproduced with permission from Nat. Photonics 7, 378-381 (2013). Copyright 2013 Springer Nature. (b) D. Huang, Sci. Rep., 6, 1-9, 2016; licensed under a Creative Commons Attribution ($C$ BY) license. (c) Reproduced with permission from Phys. Rev. Lett. 125, 010502 (2020). Copyright 2020 American Physical Society.
@@
The key modules of an in-line LO CV-QKD system are the multiplexing module and the reconciliation module. It can be seen that the development of the reconciliation directly supports the long-distance transmission, where the long transmission distance always combines with high reconciliation efficiency. The multiplexing scheme finally evolves into the time and polarization multiplexing, since this is the simplest method to achieve sufficient isolation between quantum signal and LO.
Figure 27: The field tests of in-line LO CV-QKD systems. (a) The first field test of CV-QKD system, from S. Fossier et al. Fossier et al. (2009b). (b) The field test of using CV-QKD for classical symmetric encryption, from P. Jouguet et al. Jouguet et al. (2012a) ©2012 Google. (c) Field demonstration of a CV-QKD network, from D. Huang et al. Huang et al. (2016a). (d) The longest field test of CV-QKD with 50 km commercial fiber, from Y. Zhang et al. Zhang et al. (2019a). (a) Reproduced with permission from New J. Phys. 11, 045023 (2009). Copyright 2009 IOP Publishing Ltd. (b) Copyright 2012 Google. (c) Copyright 2016 Google. (d) Reproduced with permission from Quantum Sci. Technol. 4, 035006 (2019). Copyright 2019 IOP Publishing Ltd.
@@
Besides the reconciliation and multiplexing, the modulation, monitoring, sampling and compensation techniques are also well developed. The generation of light pulses with high extinction ratio is demonstrated in 2015 Wang et al. (2015c), where the extinction ratio overpasses 80 dB by using a double cascaded MZI modulation. Later, the imperfect quantum state preparation is theoretically and experimentally investigated in 2017, which demonstrates that the incorrect calibration of the working parameters for the amplitude modulator and phase modulator can lead to a significant increase of the excess noise and misestimate of the channel loss. Schemes for calibrating the working parameters of the modulators are proposed and demonstrated to solve the imperfect state preparation issue. An LO monitoring scheme is proposed to enhance the practical security of a CV-QKD system Liu et al. (2017b). The LO is monitored by the balanced photodiode, and the SNU is real-time calibrated. The excess noise in the system and its impact on the performance is also investigated Wang et al. (2019d). Finite sampling bandwidth of the analog-to-digital (AD) converter may lead to inaccurate results of pulse peak sampling, which is solved by a dynamic delay adjusting module and a statistical power feedback-control algorithm Li et al. (2016d).